# Opt-Out: Investigating Entity-Level Unlearning for Large Language Models via Optimal Transport

### v2

@@ Abstract
Minseok Choi       Daniel Rim       Dohyun Lee       Jaegul Choo KAIST AI     Hyundai Motor Company {minseok.choi,aiclaudev,jchoo}@kaist.ac.kr drim@hyundai.com
@@ 1 Introduction
Figure 1: Motivation for entity-level unlearning. When a user submits an unlearning request, the goal of entity-level unlearning is to remove all information related to the specified entity while carefully preserving knowledge about neighboring entities and the broader world knowledge possessed by the LLM.
@@
In this work, we investigate entity-level unlearning, which focuses on removing all knowledge associated with a specific entity while retaining the rest of the model’s information. To simulate real-world unlearning scenarios, we introduce the first Entity-Level Unlearning Dataset (ELUDe), consisting of 20 real-world target entities built from their respective Wikipedia pages. Additionally, we create a synthetic dataset of 10 neighboring entities for each target, serving as retention data that is closely related to the target entity but should remain unforgotten. Our findings indicate that incorporating neighboring knowledge during training is critical for LLMs to better distinguish what to forget and retain. To further improve the performance of entity-level unlearning, we propose Opt-Out, a novel fine-grained unlearning method grounded in optimal transport theory. Specifically, Opt-Out employs the Wasserstein distance from the LLM’s initial weights to regularize the unlearning process with the optimal transportation cost between the parameters. This enables fine-grained control over the parameters, maximizing those crucial for unlearning while minimizing those essential for retention. We evaluate our framework on ELUDe, alongside several LLM benchmarks, and demonstrate that Opt-Out outperforms existing unlearning methods in both unlearning and retaining performance, highlighting the effectiveness of our approach.
@@ 2.2 Selecting Neighboring Entities
We can employ any kind of textual data for retention, such as TruthfulQA Yao et al. (2023) or Wikitext Li et al. (2024). However, we posit that finetuning with a general retain set may not effectively disentangle entity knowledge from related information. Inspired by hard negatives in representation learning Gillick et al. (2019), we curate the retained data by mining neighboring pages of a given entity. Specifically, for each target entity, we select 10 neighboring entities based on the following criteria: 1) there is a bidirectional relationship, meaning both entities link to each other and are mentioned at least once on their respective pages, 2) the neighboring pages rank within the top 10 in terms of page views over the past three years, and 3) the neighboring pages are all people. For 20 target entities, this process yields 200 neighboring entities; however, due to overlap, we identify 144 unique neighboring entities.
@@ 2.3 Generating QA Pairs
After identifying all entities, we transform each corresponding Wikipedia page into a set of QA pairs. While it is technically feasible to input the entire page into an LLM, we find that the QA format works more seamlessly with chat-based models and better simulates real-world interactions. To generate these QA pairs, we process each paragraph through GPT-4o Achiam et al. (2023), prompting it to create as many QA pairs as possible, aiming to cover the full scope of factual content. The specific prompt used for this data generation process is detailed in Appendix D. Since some paragraphs may convey overlapping or identical information, we apply a deduplication step using BERT embeddings from Sentence Transformer Reimers and Gurevych (2019). On average, approximately 647 QA pairs per entity were created.
@@ 3.1 Problem Definition
Given a token sequence $\mathbf{x}=\{x\}_{i=1}^{T}$ within the training dataset $\mathcal{D}=\{\mathbf{x}\}_{i=1}^{N}$, the goal of knowledge unlearning is to safely eliminate the influence of a specific subset of data $\mathcal{D}_{f}$ from a trained machine learning model. This process ensures that the model behaves as though the removed data was never included in the training while preserving its performance on the remaining dataset. Conventionally, the data to be forgotten $\mathcal{D}_{f}$ is referred to as the forget set, and the data to be retained is called the retain set. For simplicity, we focus on the standard scenario where $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$ are mutually exclusive (i.e., $\mathcal{D}_{f}\cap\mathcal{D}_{r}=\varnothing$). In entity-level unlearning, we consider $\mathcal{D}_{f}^{t}$, which includes all data points related to a specific target entity $t$, while $\mathcal{D}_{r}^{t}$ consists of the remaining data that does not pertain to that entity. The objective of entity-level unlearning is to train the model $\phi_{\theta}$ so that the updated model $\phi_{\theta^{\prime}}=S(\phi_{\theta};\mathcal{D}_{f}^{t})$ reflects the removal of $\mathcal{D}_{f}^{t}$. The unlearning function $S$ ensures the model operates as if it was trained exclusively on $\mathcal{D}_{r}^{t}$, forgetting $\mathcal{D}_{f}^{t}$ while retaining its performance on $\mathcal{D}_{r}^{t}$.
@@ 3.2 Knowledge Unlearning
where $\sigma$ represents the sigmoid function, $\eta>0$ is the inverse temperature, and $\phi_{\text{ref}}$ is a reference model. In entity-level unlearning, we observe that the NPO loss also produces much more stable and reliable results in practice. However, finetuning solely on the forget set eventually leads to model degradation and collapse. As with prior unlearning methods, we also train the model on the retain set to explicitly preserve the remaining knowledge. This is achieved through standard language modeling on the retain set, which serves as the positive counterpart to Equation 1:
$\mathcal{L}_{\text{RT}}=-\mathbb{E}_{\mathcal{D}_{r}^{t}}[\log(\phi_{\theta}(y|x))].$ (3)
@@ 4.2 Evaluation Metrics
For Forget Quality (FQ), we compute the harmonic mean of the three values on the forget set [^3] Unlike in Maini et al. (2024), we do not use the $p$-value from the Kolmogorov-Smirnov test as FQ because it is impossible to compare against a perfectly unlearned model in our setup, and even more so in real-world applications., while for Retain Quality (RQ), we take the harmonic mean of the six values across both the retain and world sets to prevent low scores from getting averaged out. Some values are inverted so that higher values indicate better performance (e.g., $\text{max}(0,1-R_{\text{truth}})$ is used in RQ).
Figure 3: Forget Quality performance (%) of different +RT methods following unlearning on Llama-3.1-8B-Instruct, evaluated against nine types of adversarial prompt attacks. Each attack is described in detail in Appendix E.
@@ 4.5 Performance Against LLM Attacks
Given the use of instruction-tuned models, safeguarding against malicious prompt attacks is vital. To rigorously evaluate the efficacy of unlearning in mitigating adversarial attacks, we follow Jin et al. (2024) and assess unlearned models against nine different types of adversarial threats. Detailed descriptions of the attack examples are provided in Appendix E. As illustrated in Figure 3, our proposed approach, Opt-Out, consistently achieves high-quality forgetting across various adversarial attacks, demonstrating strong robustness against malicious prompts.
Distance Metric FQ RQ Util. Wasserstein (ours) 87.8 46.6 63.3 Manhattan 47.0 50.9 64.6 Euclidean 81.5 46.2 63.0 Chebyshev 86.3 45.4 62.2 1 - Cosine Similarity 81.6 45.8 62.8
@@ 6 Conclusion
While our framework shows promising performance in unlearning entity-level knowledge, several areas warrant further development. First, although our approach significantly improves Forget Quality, it still results in a slight performance decline when it comes to retaining other knowledge. This is evident in Retain Quality and model utility scores compared to the original model, highlighting a valuable opportunity for future research on selectively forgetting while maintaining overall model performance. Second, our method remains susceptible to generating gibberish post-unlearning. Although it effectively removes parametric knowledge, ensuring the LLM functions correctly for a seamless end-user experience in real-world deployment remains an issue. Combining with the IDK method or remapping outputs to automated responses after unlearning could be considered for a simple fix. Lastly, due to computational constraints, we were unable to test models at the scale of 70B parameters or larger. Exploring unlearning techniques with much larger models would better align with the behavior of proprietary models.
@@ Appendix B Human Evaluation
Forget Set Retain Set Relevance 85.7 98.3 Diversity 88.7 99.0 Factuality 99.7 100.0
@@
Appendix C Full Evaluation Results
[…]
@@ Appendix C Full Evaluation Results
Appendix D Prompts
@@
Appendix E Dataset Examples

### v3

@@ Abstract
Minseok Choi       Daniel Rim†, ‡       Dohyun Lee       Jaegul Choo KAIST AI     Hyundai Motor Company {minseok.choi,aiclaudev,jchoo}@kaist.ac.kr drim@hyundai.com
@@ 1 Introduction
Figure 1: Motivation for entity-level unlearning. While instance-level unlearning operates at a coarse granularity, entity-level unlearning aims to remove target entity knowledge while carefully preserving knowledge about neighboring entities and the broader world knowledge.
@@
In this work, we investigate entity-level unlearning, which focuses on removing all knowledge associated with a specific entity while retaining the rest of the model’s information. To simulate real-world unlearning scenarios, we introduce the first Entity-Level Unlearning Dataset (ELUDe), consisting of 20 real-world target entities built from their respective Wikipedia pages. Additionally, we create a dataset of 10 neighboring entities for each target, serving as retention data that is closely related to the target entity but should remain unforgotten. To further improve the performance of entity-level unlearning, we propose Opt-Out, a novel fine-grained unlearning method grounded in optimal transport theory. Specifically, Opt-Out employs the Wasserstein distance from the LLM’s initial weights to regularize the unlearning process with the optimal transportation cost between the parameters. This enables fine-grained control over the parameters, maximizing those crucial for unlearning while minimizing those essential for retention. We evaluate our framework on ELUDe, alongside several LLM benchmarks, and demonstrate that Opt-Out outperforms existing unlearning methods in both unlearning and retaining performance, highlighting the effectiveness of our approach. Our work focuses on Wikipedia entities due to their extensive coverage and accessibility, rather than the actual privacy data; however, we hope this work provides a testbed for entity-level unlearning, taking a modest step toward advancing the development of practical unlearning methods.
@@ 2.2 Selecting Neighboring Entities
We can employ any kind of textual data for retention, such as TruthfulQA Yao et al. (2023) or Wikitext Li et al. (2024). However, we posit that finetuning with a general retain set may not effectively disentangle entity knowledge from related information. Inspired by hard negatives in representation learning Gillick et al. (2019), we curate the retained data by mining neighboring pages of a given entity. Specifically, for each target entity, we select 10 neighboring entities based on the following criteria: 1) there is a bidirectional relationship, meaning both entities link to each other and are mentioned at least once on their respective pages, 2) the neighboring pages rank within the top 10 in terms of page views over the past three years, and 3) the neighboring pages are all people. For 20 target entities, this process yields 144 unique neighboring entities (due to overlap).
@@ 2.3 Generating QA Pairs
After identifying all entities, we transform each corresponding Wikipedia page into a set of QA pairs. While it is technically feasible to input the entire page into an LLM, we find that the QA format works more seamlessly with chat-based models and better simulates real-world interactions. To generate these QA pairs, we process each paragraph through GPT-4o Achiam et al. (2023), prompting it to create as many QA pairs as possible, aiming to cover the full scope of factual content. The specific prompt used for this data generation process is detailed in Appendix F. Since some paragraphs may convey overlapping or identical information, we apply a deduplication step using BERT embeddings from Sentence Transformer Reimers and Gurevych (2019). On average, approximately 647 QA pairs per entity were created.
@@ 3.1 Problem Definition
Knowledge Unlearning
Given a token sequence $\mathbf{x}=\{x\}_{i=1}^{T}$ within the training dataset $\mathcal{D}=\{\mathbf{x}\}_{i=1}^{N}$, the goal of knowledge unlearning is to safely eliminate the influence of a specific subset of data $\mathcal{D}_{f}$ from a trained machine learning model. This process ensures that the model behaves as though the removed data was never included in the training while preserving its performance on the remaining dataset. Conventionally, the data to be forgotten $\mathcal{D}_{f}$ is referred to as the forget set, and the data to be retained is called the retain set. For simplicity, we focus on the standard scenario where $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$ are mutually exclusive (i.e., $\mathcal{D}_{f}\cap\mathcal{D}_{r}=\varnothing$).
Entity-Level Unlearning
Entity-level unlearning aims to remove all knowledge related to a specific target entity from a trained machine learning model. However, because the knowledge about an entity within a model is inherently abstract and cannot be explicitly predefined, we adopt the framework proposed by Ma et al. (2024). This reformulates the task as erasing all knowledge associated with a target entity by identifying and deleting a specific subset of entity-associated knowledge, which we term the “proxy forget set”. The proxy forget set encapsulates identifiable aspects of the target entity’s knowledge. Since it is equally impractical to obtain the exact retain set used during pretraining, we define the “proxy retain set” as a specific subset of knowledge encompassing everything except the target entity.
Formally, for a target entity $t$ and its proxy forget set $\mathcal{D}_{f}^{t}$, the objective of entity-level unlearning is to train a model $\phi_{\theta}$ such that the updated model $\phi_{\theta^{\prime}}=S(\phi_{\theta};\mathcal{D}_{f}^{t})$ operates as if it were trained exclusively on the proxy retain set $\mathcal{D}_{r}^{t}$. The unlearning function $S$ ensures that $\mathcal{D}_{f}^{t}$ is effectively forgotten while retaining the model’s performance on $\mathcal{D}_{r}^{t}$. It is important to note that the exact forget and retain sets are unattainable, except in highly controlled scenarios such as fictitious unlearning Maini et al. (2024). Consequently, achieving high knowledge coverage in the proxy sets directly improves alignment with the ideal performance attainable using exact sets. To address this, we leverage Wikipedia as our knowledge source due to its extensive coverage of entity-specific information, thereby narrowing the gap between proxy and exact sets.
@@ 3.2 Knowledge Unlearning
where $\sigma$ represents the sigmoid function, $\eta>0$ is the inverse temperature, and $\phi_{\text{ref}}$ is a reference model. In entity-level unlearning, we observe that the NPO loss also produces much more stable and reliable results in practice. However, finetuning solely on the forget set eventually leads to model degradation and collapse. As with prior unlearning methods, we also train the model on the retain set to explicitly preserve the remaining knowledge. This is achieved through standard language modeling on the retain set, which serves as the positive counterpart to Equation 1: $\mathcal{L}_{\text{RT}}=-\mathbb{E}_{\mathcal{D}_{r}^{t}}[\log(\phi_{\theta}(y|x))]$.
@@ 4.2 Evaluation Metrics
For Forget Quality (FQ), we compute the harmonic mean of the three values on the forget set [^3] Unlike in Maini et al. (2024), we do not use the $p$-value from the Kolmogorov-Smirnov test as FQ because it is impossible to compare against a perfectly unlearned model in our setup, and even more so in real-world applications. For comparison with exact unlearning, refer to Appendix E for our results on the TOFU dataset., while for Retain Quality (RQ), we take the harmonic mean of the six values across both the retain and world sets to prevent low scores from getting averaged out. Some values are inverted so that higher values indicate better performance (e.g., $\text{max}(0,1-R_{\text{truth}})$ is used in RQ).
Figure 3: Forget Quality performance (%) of different +RT methods following unlearning on Llama-3.1-8B-Instruct, evaluated against nine types of adversarial prompt attacks. Each attack is described in detail in Appendix G.
@@ 4.5 Performance Against LLM Attacks
Llama-3.1 Phi-3.5
Method mean std mean std
Oracle 50.0 - 50.0 -
GA 53.8 2.9 54.7 1.4
DPO 56.2 3.8 53.8 2.2
NPO 53.8 1.8 52.8 1.6
IDK 59.3 2.3 58.7 1.8
GA+RT 50.9 3.1 49.7 1.0
DPO+RT 50.3 3.1 49.7 3.1
NPO+RT 49.6 2.5 50.9 2.5
IDK+RT 69.1 1.5 67.3 1.9
Opt-Out (ours) 48.6 1.0 49.1 1.1
Table 2: MIA accuracy (%) of a trained binary classifier (“the attacker”) predicting whether an input data belongs to the training set. 50% indicates the best performance.
@@
Given the use of instruction-tuned models, safeguarding against malicious prompt attacks is vital. To rigorously evaluate the efficacy of unlearning in mitigating adversarial attacks, we follow Jin et al. (2024) and assess unlearned models against nine different types of adversarial threats. Detailed descriptions of the attack examples are provided in Appendix G. As illustrated in Figure 3, our proposed approach, Opt-Out, consistently achieves high-quality forgetting across various adversarial attacks, demonstrating strong robustness against malicious prompts.
@@ 4.6 Effect of Wasserstein Regularization
Distance Metric FQ RQ Util. Wasserstein (ours) 87.8 46.6 63.3 Manhattan 47.0 50.9 64.6 Euclidean 81.5 46.2 63.0 Chebyshev 86.3 45.4 62.2 Cosine 81.6 45.8 62.8
@@ 6 Conclusion
While our framework shows promising performance in unlearning entity-level knowledge, several areas warrant further refinement. First, our work focuses on unlearning Wikipedia entities, which may differ slightly from erasing data related to actual users. Nevertheless, creating meaningful forget and retain sets for an arbitrary person (e.g., Alice) is challenging, as it is difficult to capture how much the LLM knows about her. Therefore, we have leveraged Wikipedia, where the pages themselves serve as a useful proxy for comprehensive data coverage of a particular entity, enabling effective evaluation of full entity-level erasure. Future work could extend our approach to real-world privacy data, incorporating advanced anonymization techniques to better align with practical use cases. Second, our method remains susceptible to generating gibberish post-unlearning. Although it effectively removes parametric knowledge, ensuring the LLM functions correctly for a seamless end-user experience in real-world deployment remains an issue. Combining with the IDK method or remapping outputs to automated responses after unlearning could be considered a simple fix. Lastly, due to computational constraints, we were unable to test models at the scale of 70B parameters or larger. Exploring unlearning techniques with much larger models would better align with the behavior of proprietary models.
Acknowledgements
We are deeply grateful to Hojoon Lee and Dongyoon Hwang for generously sharing their expertise during our discussions. We also appreciate the thoughtful feedback provided by the anonymous reviewers. This work was supported by the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST), RS-2025-02304967, AI Star Fellowship (KAIST)), and Samsung Electronics Co., Ltd.
@@ Appendix B Human Evaluation
Forget Set Retain Set Relevance 89.0 96.0 Diversity 90.0 99.5 Factuality 99.0 98.5
@@
Appendix C Computational Cost Analysis
To portray the feasibility of Opt-Out, we conduct a computational cost analysis and present the results in Table 5. Opt-Out introduces a small additional overhead during training while maintaining an identical inference cost to existing methods, ensuring its practicality for large-scale applications. Specifically, the additional overhead stems from computing the Sliced Wasserstein Distance (SWD) between two sets of parametric weights. Regarding memory usage during training, Opt-Out is comparable to DPO and NPO, which also require holding an extra reference (frozen) model. The additional memory cost in Opt-Out arises from a lightweight random projection layer used in SWD computation.
Given that $L$ is the input sequence length, the training time and memory complexities of Transformer is $O(L^{2})$. +RT incurs additional overhead due to an extra forward pass, while DPO/NPO require a third forward pass through a reference model (DPO demands extra two passes for the preference loss). Opt-Out adds a small training overhead of $O(L\log L)$ for SWD computation (due to sorting) and $O(Ld)$ for projected representations, where $d$ is the dimension of the random projection layer and is kept small. We believe these additional costs are manageable even at scale, as the $L\log L$ term grows significantly slower compared to the $L^{2}$ complexity of Transformer training itself.
Training Inference
Method Time Memory Time Memory
GA+RT $\mathcal{O}(2L^{2})$ $\mathcal{O}(L^{2})$ $\mathcal{O}(L)$ $\mathcal{O}(L^{2})$
DPO+RT $\mathcal{O}(5L^{2})$ $\mathcal{O}(2L^{2})$ $\mathcal{O}(L)$ $\mathcal{O}(L^{2})$
NPO+RT $\mathcal{O}(3L^{2})$ $\mathcal{O}(2L^{2})$ $\mathcal{O}(L)$ $\mathcal{O}(L^{2})$
IDK+RT $\mathcal{O}(2L^{2})$ $\mathcal{O}(L^{2})$ $\mathcal{O}(L)$ $\mathcal{O}(L^{2})$
Opt-Out $\mathcal{O}(3L^{2}+L\log L)$ $\mathcal{O}(2L^{2}+Ld)$ $\mathcal{O}(L)$ $\mathcal{O}(L^{2})$
Table 5: Computational complexity of various methods (constants included to better depict the comparison). Opt-Out incurs a slight additional overhead during training while maintaining an identical inference cost to existing methods, ensuring its practicality.
@@ Appendix C Computational Cost Analysis
Appendix D Full Evaluation Results
[…]
Appendix E Exact Unlearning Results
We conduct additional experiments on the TOFU dataset Maini et al. (2024) to further demonstrate the effectiveness of our method and report the results in Table 7. The TOFU dataset is entirely fictitious and was not included during the model’s pretraining. This ensures that the model starts as “perfectly unlearned”, enabling a direct comparison between our approach and exact unlearning. Forget Quality (FQ) in this context is calculated using $p$-values derived from the Kolmogorov-Smirnov test (KS-Test), which quantifies the divergence between the distributions of Truth Ratios in unlearned models and the exact unlearning model. Retain Quality (RQ) is also slightly adapted here, reflecting the harmonic mean of nine values (Probability, ROUGE, and Truth Ratio measured across Retain Set, Real Authors, and World Facts). For clarity, an asterisk (*) differentiates terms redefined in this context from those in our main paper. The results demonstrate that Opt-Out consistently outperforms other approaches, highlighting its robustness.
Appendix F Prompts
@@
Llama-3.1 Phi-3.5
Method FQ* RQ* FQ* RQ*
GA+RT 0.0 16.6 46.6 45.2
DPO+RT 0.0 47.7 1.6 44.7
NPO+RT 11.2 46.1 71.3 45.9
IDK+RT 0.0 41.0 22.1 39.3
Opt-Out (ours) 86.6 40.6 86.6 45.9
Table 7: Performance (%) of various methods after unlearning fictitious data in TOFU (5% forget set setting). FQ* is calculated using $p$-values derived from the KS-Test, and RQ* reflects the harmonic mean of nine values (three metrics measured across three subsets in TOFU).
Appendix G Dataset Examples